40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 18:53:48 2013, Jenne, Update, Alignment, IFO ready for doors, then pumping 
    Reply  Wed Feb 20 19:48:16 2013, yuta, Update, Alignment, BS table oplev re-arranged IFOALIGN_QPDs_OSEMs.png
       Reply  Wed Feb 20 19:58:59 2013, rana, Update, Alignment, BS table oplev re-arranged 
          Reply  Wed Feb 20 20:15:29 2013, yuta, Update, Alignment, SRM oplev status 
Message ID: 8117     Entry time: Wed Feb 20 18:53:48 2013     Reply to this: 8119
Author: Jenne 
Type: Update 
Category: Alignment 
Subject: IFO ready for doors, then pumping 

[Yuta, Manasa, Sendhil, Jenne, Steve, Jamie, Koji, Evan]

The interferometer is well-aligned, and ready for pump-down.  The access connector is in place, as are the ETM heavy doors.  We will do ITM and BS doors tomorrow, then begin pumping.

Before we redid the ITM pointing, I confirmed that I could see both POX and POY on their respective tables, on a camera, unclipped.  I should check again quickly now that the ITM pointing has been finalized.

We went back to the arms, to perfect the ITM pointing.  Input beam was already centered at ETMY.  ETMY was pointing so that beam reflected to ITMY.  ITMY was adjusted a few (less than 4?) steps of 1e-3 size, to make reflected beam hit center of ETMY. 

BS was already pointing so beam hit center of ETMX.  ETMX was pointing to hit center of ITMX.  ITMX was adjusted a few (less than 4 again?) steps of 1e-3 size to make reflected beam hit center of ETMX.

Checked centering on AS path.  AS beam comes out of the vacuum a little low, but this wasn't discovered until after the access connector was in place.  We could adjust PZT3 (last AS mirror on BS table that sends beam over to OMC table), but we don't want to do this since we won't be able to re-confirm centering on the 3 mirrors on the OMC table.

Green beams (first Y, then X) were aligned using out-of-vac steering mirrors until beams were flashing in their respective arm cavities.  Green Y is a little close to the edge of the bottom periscope mirror, on the "up" periscope.  Since there is no steering between the arm and this periscope, fixing it would require moving the periscope.  We leave this to the next vent, when we finally install the BS table extension.  We were flashing a higher order yaw mode (5ish nodes) for the Y arm, and the very edge of the higher order mode on one side was a little bit clipped after reflecting off the steering mirror on the OMC table.  This is happening because that mirror is in the mount backwards (so we have access to the knobs).  We are confident that the straight-through beam is well centered on that mirror, so once we get it aligned to TEM00, there will be no clipping. We then did the X arm green, which was flashing a pitch higher order mode (again 5ish nodes).  The very edge of the higher order mode is clipping a little bit on the top mirror of the "down" periscope on the IMC table, but again the straight through beam is okay, and we're confident that the TEM00 mode will make it unclipped.  We could have touched some steering mirrors on the BS table, but since this was once upon a time well aligned, we don't want to futz with it.

Corner oplevs are all centered on their QPDs.  (The ETM oplevs were centered a few days ago).

Access connector and ETM doors are on.

The last 3 vertex doors will go on tomorrow when Steve gets in, and then we'll start pumping. 

There are no in-vac PZTs that need to be turned on (we've been using the output steering PZTs as non-energized fixed mirrors for some time), so we can lock at our leisure tomorrow afternoon.

ELOG V3.1.3-